### v1

# Asymmetric Symmetry Breaking: Unequal Probabilities of Vacuum Selection

Abstract
We study the probabilities of a field, subject to random perturbations, to roll down from the top of a potential, where the top is only $C^{1}$ continuous. We find that the probability to roll down to the left or right depends the square root of the second derivative of the potential at the top. We solve this problem theoretically by using the Fokker-Planck equations in stochastic process and verify our findings numerically. This study may potentially be a new mechanism to explain the origins of asymmetries in the Universe.
I Introduction
Spontaneous symmetry breaking (SSB) is an important concept in contemporary physics. It is vital for the superconducting phenomena in condensed matter physics [1], the Higgs mechanism in particle physics [2] and the cosmological phase transitions in cosmology [3], etc. A typical picture of SSB is a field to roll down a potential with Mexican-hat profile. The field was sitting at the top of the potential initially with higher symmetries. Since the top is unstable, the field will then roll down the potential and settle down in one of the lowest energy states. Compared to the symmetries at the top, the chosen state has lower symmetries, i.e., symmetry was broken. We call this process of SSB as ‘vacuum selection’.
Commonly, the probabilities for the field to roll down the potential are identical, that is to say, the probabilities of the vacuum selection are equal. For instance, if the potential is only one-dimensional in the field, then the filed to roll down from the top to the left or right are exactly one-half. A natural question is: how can we have unequal probabilities for the vacuum selection. If we could realize it, it may have important meanings to the physics of symmetry breaking, such as the cosmological phase transitions we mentioned above.
In this work, we utilize a special Higgs-type potential to realize the unequal probabilities of the vacuum selection. In particular, the top of the potential is only $C^{1}$ continuous, but discontinuous at higher order, see the sketchy pictures in Fig.1. We can call this potential as $\mathring{C}^{1}$ continuous, in which the upper circle indicates that only the top is $C^{1}$ continuous while other parts are $C^{\infty}$ continuous.
Figure 1: Schematic picture to illustrate the rolling down of the fields. The system is subject to thermal perturbations, and the top of the potential is only $C^{1}$ continuous. Fields tend to roll to the steeper side.
The system is subject to random external forces, for example coupled to a thermal reservoir. Under the random perturbations, the field at the top will have the probability to roll down to the left or right side. And eventually settle down to the lower energy states at the minimum of the potential. We use the Fokker-Planck equation in stochastic process to solve this problem theoretically, and find that the probability for the field to roll down to the left (right) depends on the square root of the second derivative of the potential at left (right) side of the top. The we adopt the time-dependent Ginzburg-Landau model to simulate this process by virtue of two kinds of potentials: one is the Higgs-type potential, the other is the cosine-type potential. In both models, we see that the numerical results are consistent with the theoretical predictions very well.
The unequal probabilities of the vacuum selection is key to the fundamental structures of the vacuum. Therefore, our work may have significant implications to the phase transitions induced by spontaneous symmetry breaking.
II Theory
We consider the time-dependent Ginzburg-Landau equation with an external random force $F(t)$,
$\frac{d\phi}{dt}+\Gamma\frac{\partial V(\phi)}{\partial\phi}=F(t)$ (1)
where $\phi$ is a scalar field and $\Gamma$ is a positive coefficient, $F(t)$ is the external random force satisfying the Gaussian white noise $\langle F(t)\rangle=0,\langle F(t)F(t^{\prime})\rangle=Q\delta(t-t^{\prime})$ where $Q$ is the fluctuation strength. If the random perturbation comes from the thermal reservoir, $Q$ will be linearly proportional to the temperature $T$. From the stochastic process, the probability to find the field in the range $[\phi,\phi+d\phi]$ at instant $t$ is $P(\phi,t)$, which satisfies the Fokker-Planck equation [4],
$\frac{\partial P}{\partial t}=\Gamma\frac{\partial}{\partial\phi}\left(\frac{\partial V}{\partial\phi}P\right)+\frac{Q}{2}\frac{\partial^{2}P}{\partial\phi^{2}}.$ (2)
Assume that the potential behaves like a Mexican hat in Fig.1, and at initial time the field is sitting at the top of the potential. Without loss of generality, we can always translate the top of the potential to be seated at $\phi=0$. Hence, we can Taylor expand the potential near the top and the Fokker-Planck equation (2) reduces to
$\frac{\partial P}{\partial t}=\Gamma V^{\prime\prime}_{0}\frac{\partial}{\partial\phi}\left(\phi P\right)+\frac{Q}{2}\frac{\partial^{2}P}{\partial\phi^{2}}$ (3)
where $V^{\prime\prime}_{0}$ indicates $\partial^{2}V(\phi)/\partial\phi^{2}|_{\phi\sim 0}$. In the Eq.(3) the term containing $V^{\prime}$ vanishes since the top is a maximum and $C^{1}$ continuous. Higher terms of the expansion will make minor contributions compared to the linear term of $\phi$ near $\phi\sim 0$. The stationary solution of the above equation, i.e, $\partial P/\partial t=0$, is
$P_{\rm st.}(\phi)=\sqrt{\frac{\Gamma V^{\prime\prime}_{0}}{\pi Q}}e^{-\frac{\Gamma V^{\prime\prime}_{0}\phi^{2}}{Q}}$ (4)
which is a normal distribution centered at $\phi=0$.
If at the initial time $\phi$ sits at the top of the potential, i.e., $\phi=0$. Then we evolve the system to see whether $\phi$ will roll down to the left or right. Random fluctuations of $F(t)$ in this case plays a role of disturbing $\phi$ from $\phi=0$ to $\phi=\epsilon$, where $\epsilon$ is a relatively small value. If $\epsilon$ is negative (positive), the field $\phi$ will roll left (right) continuously. Since $\epsilon$ is a small value, then we can ignore the exponential term in the Eq.(4). Therefore, the probability to roll down to the left/right directly depends the factor $\sqrt{{\Gamma V^{\prime\prime}_{0}}/{(\pi Q)}}$. Now if the potential is only $\mathring{C}^{1}$ continuous, the second derivative of the potential $V^{\prime\prime}_{0}$ near $\phi\sim 0$ are different. If we define $V^{\prime\prime}_{\pm}=\partial^{2}V/\partial\phi^{2}$ at $\phi=0_{\pm}$, we then get
$\frac{P_{+}}{P_{-}}=\frac{\sqrt{{\Gamma V^{\prime\prime}_{+}}/{(\pi Q)}}}{\sqrt{{\Gamma V^{\prime\prime}_{-}}/{(\pi Q)}}}=\sqrt{\frac{V^{\prime\prime}_{+}}{V^{\prime\prime}_{-}}}.$ (5)
where $P_{-}(P_{+})$ indicates the probability to roll down left (right). Therefore, we see that the probability to roll down left or right depends on the square root of the second derivatives of the potential at the top. Since $P_{+}+P_{-}=1$ and from Eq.(5), it is readily to see that
$P_{+}=\frac{\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}}{1+\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}}.$ (6)
Or, equivalently, $P_{-}=\frac{\sqrt{V^{\prime\prime}_{-}/V^{\prime\prime}_{+}}}{1+\sqrt{V^{\prime\prime}_{-}/V^{\prime\prime}_{+}}}$.
III Numerical Evidences
We will use two different potentials to verify our conclusions, i.e., Eq.(5) and (6) above. Specifically, the first potential is the Higgs-type potential, and the second one is the cosine-type potential.
III.1 Higgs-type potential
aRefer to captionb
Refer to captioncRefer to captiond
Figure 2: a. The Higgs-type potential $V(\phi)$ (blue) in Eq.(7) with $\alpha_{-}=1$ and $\alpha_{+}=2$, and its derivative $V^{\prime}(\phi)$ (red); b. Time time evolution of the scalar fields. Each has three cases to approach positive and negative minimums. The inset plot shows the very early fluctuations around the top due to the external temperature. After some time, they will persistently roll down the potential as the arrows indicate; c. Relation between the probability $P_{+}$ to the ratio $\alpha_{+}/\alpha_{-}$. Numerical results match the theoretical prediction (red line) Eq.(6) very well; d. Relations between the ratio $P_{+}/P_{-}$ to the ratio $\alpha_{+}/\alpha_{-}$. The numerical results are completely consistent with the theoretical prediction (red line) in Eq.(5). The inset plot shows the double logarithmic plot of $P_{+}/P_{-}$ against $\alpha_{+}/\alpha_{-}$, in which the slope $1/2$ exactly means the square root relations in Eq.(5).
The Higgs-type potentials is
$V(\phi)$ $=$ $\theta(-\phi)(-\alpha_{-}\phi^{2}+\phi^{4})$ (7)
$+\theta(\phi)(-\alpha_{+}\phi^{2}+\phi^{4}),$
where $\theta$ is the Heaviside step function with $\theta(x>0)=1,\theta(x<0)=0$ and $\theta(0)=1/2$. In numerics, we fix $\alpha_{-}\equiv 1$ and vary $\alpha_{+}$ in the range $\alpha_{+}\in[0.05,5]$. Please find the profile of the potential and its derivative in panel a of Fig.2, in which we set $\alpha_{+}=2$. Obviously, the profile of the potential satisfies our requirement of $\mathring{C}^{1}$ continuous. In the numerics, we use the 4th Runge-Kutta methods to evolve the system. And we have set the time step as $\Delta t=0.1$, the coefficient $\Gamma=5$ and the temperature $T=0.01$. For the statistics, we independently simulate the dynamics for 30000 times.
In the panel b of Fig.2 we show the time evolution of the scalar fields with the potential in panel a. In this plot we show three cases for the scalar field to select the positive vacuum and three cases for selecting the negative vacuum. Because of the random temperature perturbations, each three lines are not overlap. But at the final time they will overlap and reach the equilibrium values of the positive/negative vacuum, respectively. The positive (negative) equilibrium value is $1$ ($-0.707$) which are exactly the value of $\phi_{+}=1$ ($\phi_{-}=-1/\sqrt{2}$) corresponding to the extremal minimums of the potential. In the inset plot of panel b we show the early evolutions of two cases of the fields. We can see that in the very early time the fields will fluctuate around $\phi=0$ due to the external temperature. But this fluctuation near the top of the potential will not last long enough because they are unstable at the top. At certain time they will go away from the top, i.e. $\phi=0$, and then persistently roll down the hill until meeting the minimum of the potential. The arrows in the inset plot imply the persistence of rolling down the hill. Finally, they will settle down at the minimums of the potential, which is reflected by the plateaus in panel b.
Panel c of Fig.2 show the relation between $P_{+}$ and $\alpha_{+}/\alpha_{-}$. From the potential (7) we see that $\alpha_{+}/\alpha_{-}={V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$ at the top of the potential. From panel c we see that the numerical results match the theoretical prediction $P_{+}=\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}/\left(1+\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}\right)$ in Eq.(6) very well.
Panel d of Fig.2 exhibits the relation between $P_{+}/P_{-}$ and $\alpha_{+}/\alpha_{-}$. Therefore, we see that the numerical data also match the theoretical prediction $P_{+}/P_{-}=\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$ in Eq.(5) very well. The inset plot in panel d shows the double logarithmic plots of the relation, and the slope $1/2$ exactly means the square root relation in $P_{+}/P_{-}=\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$.
III.2 Cosine-type potential
The second example to check the correctness of Eq.(5) and Eq.(6) is with the cosine-type potential,
$V(\phi)$ $=$ $\theta(-\phi)(\cos(\beta_{-}\phi)+\phi^{4})$ (8)
$+\theta(\phi)(\cos(\beta_{+}\phi)+\phi^{4}).$
In numerics, we fix $\beta_{-}\equiv 2$ and vary $\beta_{+}$ in the range $\beta_{+}\in[0.2,8]$. Now we change the time step to be $\Delta t=0.02$ and still set $\Gamma=5$ and temperature $T=0.01$.
Panel a of Fig.3 shows the profiles and the derivative of the potential (8) with $\beta_{+}=6$. Obviously, the potential satisfies our requirement of the $\mathring{C}^{1}$ continuous. The minimum values of potential sits at $\phi_{-}=-0.7937$ and $\phi_{+}=0.5089$.
Panel b of Fig.3 exhibits the time evolutions of the scalar fields. There are respectively three cases for each of them to approach $\phi_{+}$ (the upper lines) and $\phi_{-}$ (the lower lines) at equilibrium time. The inset plot shows the initial fluctuations of the scalar fields around the top, and then continuously roll down the potential as the arrows indicate.
Panel c of Fig.3 is the relation between $P_{+}$ and the ratio $\beta_{+}/\beta_{-}$. From the potential (8) we see that $\beta_{+}/\beta_{-}=\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$, therefore, the relations shown in panel c, i.e., $P_{+}=(\beta_{+}/\beta_{-})/(1+\beta_{+}/\beta_{-})$, exactly satisfies our prediction in Eq.(6).
Panel d of Fig.3 shows the relations between $P_{+}/P_{-}$ and $\beta_{+}/\beta_{-}$. The linear relation $P_{+}/P_{-}=\beta_{+}/\beta_{-}$ in numerics and theoretical predictions Eq.(5) match very well.
aRefer to captionb
Refer to captioncRefer to captiond
Figure 3: a. The cosine-type potential $V(\phi)$ (blue) in Eq.(8) with $\beta_{-}=2$ and $\beta_{+}=6$, and its derivative $V^{\prime}(\phi)$ (red); b. Time time evolution of the scalar fields. Each has three cases to approach positive and negative minimums. The inset plot shows the very early fluctuations around the top due to the external temperature. After some time, they will persistently roll down the potential as the arrows indicate; c. Relation between the probability $P_{+}$ to the ratio $\beta_{+}/\beta_{-}$. Numerical results match the theoretical prediction (red line) Eq.(6) very well; d. Relation between the ratio $P_{+}/P_{-}$ to the ratio $\beta_{+}/\beta_{-}$. The numerical results are completely consistent with the theoretical prediction (red line) in Eq.(5).
Therefore, we can see from the above two examples of different potentials that our main predictions, i.e. Eq.(5) and Eq.(6), are very accurate and match the numerical results very well. We also need to stress that the relations in Eq.(5) and Eq.(6) are robust against by changing other parameters. In numerics, we already checked that our findings in Eq.(5) and Eq.(6) are robust against: (I) varying the temperatures $T$, (II) varying the coefficient $\Gamma$, (III) adding $\phi^{3}$ terms in the potential (7) and (IV) changing the coefficients in front of $\phi^{4}$. Therefore, it reflects that our predictions Eq.(5) and Eq.(6) depending only on $\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$ are very strong and universal predictions.
IV Conclusions and Discussions
We studied the unequal probabilities of a field to roll down an asymmetric potential, in which the top is only $C^{1}$ continuous. By using the Fokker-Planck equation in stochastic equation, we theoretically found that the probability to roll down to the left (right) depends on the square root of the second derivative of the potential at that top, i.e. our main conclusions are the Eq.(5) and Eq.(6). Then we used two kinds of potentials in the time-dependent Ginzburg-Landau equation to numerically verify our conclusions. We found that our theoretical predictions are robust against varying other parameters. Therefore, our findings are universal.
There are many possible future uses for this new mechanism of asymmetric symmetry breaking. For instance, in cosmology one usually use the Higgs-type potential to study the cosmological phase transitions. If we change the potential to be like in our work, the probabilities to get the vacuum states will be different. This would change the cosmological phase transitions accordingly. Potentially, this new mechanism may play a role in the origins of asymmetries in the early Universe.
Acknowledgement
The authors thank Peng-Zhang He and Yu Zhou for the helpful discussions. HQZ would like to appreciate many interesting talks, which inspired him to think about this problem, during the conference “Workshops on Gravitation and Cosmology 2023” held in Beijing by ITP, CAS. This work was partially supported by the National Natural Science Foundation of China (Grants No.12175008).
—Appendix—
Appendix A Counting statistics of the binomial distribution
A.1 With Higgs-type potential
aRefer to captionb
Refer to captioncRefer to captiond
Figure S1: Counting statistics with Higgs-type potential in Eq.(7). a. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=2/3$; b. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=2-\sqrt{2}$; c. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2/3$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/3)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$; d. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2-\sqrt{2}$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(\sqrt{2}-1)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$.
In this appendix, we show the histograms of the probability distribution of rolling down to right. If we regard the rolling down to the right side as a success with probability $P_{+}$, then rolling down to the left is a failure with success $P_{-}=1-P_{+}$. Therefore, each time of the rolling down is a Bernoulli trial. Then after $n$ times of trials, the probability of rolling down to the right as $k$ times will satisfy the binomial distributions
$B(n,k)=\binom{n}{k}P_{+}^{k}(1-P_{+})^{n-k}.$ (S1)
In Fig.S1 we show the counting statistics of the rolling down to the right side with the Higgs-type potential in Eq.(7). For the left column, i.e., Fig.S1 a and c, we set $\alpha_{-}=1$ and $\alpha_{+}=4$. Therefore, the probability to roll down to the right side $P_{+}=\frac{\sqrt{4}}{1+\sqrt{4}}=2/3$ according to Eq.(6). While for the right column, i.e., Fig.S1 b and d, we set $\alpha_{-}=1$ and $\alpha_{+}=2$, thus $P_{+}=\frac{\sqrt{2}}{1+\sqrt{2}}=2-\sqrt{2}$. Panel a and b of Fig.S1 show the histogram of the probability $P(n,k)$, which means $k$ times of rolling down to the right side in $n$ trials. We can see that for the different probabilities $P_{+}$ and different trials numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(S1) very well.
Fig.S1 c and d shows the vanishing probability to roll down to the right in $n$ trials, i.e., $P(n,k=0)$. Theoretically, from the binomial distribution Eq.(S1), this vanishing probability should be $B(n,k=0)=(1-P_{+})^{n}$. From the two plots we see that the numerical statistics match the theoretical predictions very well.
A.2 With cosine-type potential
aRefer to captionb
Refer to captioncRefer to captiond
Figure S2: Counting statistics with cosine-type potential in Eq.(8). a. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/5$; b. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/4$; c. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/5$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(2/5)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$; d. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/4$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/4)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$.
In Fig.S2 we show the counting statistics of the rolling down to the right side with the cosine-type potential in Eq.(8). For the left column, i.e., Fig.S2 a and c, we set $\beta_{-}=2$ and $\beta_{+}=3$. Therefore, the probability to roll down to the right side $P_{+}=\frac{3/2}{1+3/2}=3/5$ according to Eq.(6). While for the right column, i.e., Fig.S2 b and d, we set $\beta_{-}=2$ and $\beta_{+}=6$, thus $P_{+}=\frac{6/2}{1+6/2}=3/4$.
From panel a and b of Fig.S2, we can see that for the different probabilities $P_{+}$ and different trials numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(S1) very well.
Fig.S2 c and d shows the vanishing probability to roll down to the right in $n$ trials, i.e., $P(n,k=0)$. Theoretically, from the binomial distribution Eq.(S1), this vanishing probability should be $B(n,k=0)=(1-P_{+})^{n}$. From the two plots we see that the numerical statistics match the theoretical predictions very well.
Therefore, the consistency between the numerics and the binomial distributions in turn implies that our predictions of the $P_{+}$ in Eq.(6) are correct.